OpenStudy (kanwal32):

Find the sum upto n terms 1+1+2/2 + 1+2+3/3 1+2+3+4/4 +.........

OpenStudy (kanwal32):

\[1+ (1+2/2) + (1+2+3/3) +( 1+2+3+4/4)\]

OpenStudy (kanwal32):

@dan815 pls help

OpenStudy (kanwal32):

@ShadowLegendX @wio @ParthKohli pls help

OpenStudy (kanwal32):

pls help

OpenStudy (triciaal):

i see sum (a + n)/n lowest 0 highest n but this is not complete need to account for the repeat

OpenStudy (kanwal32):

i have redused denominator to n(n+1)/2

OpenStudy (triciaal):

a + (a + n)/n + ...

OpenStudy (kanwal32):

but i am thinking that i should take the no of terms as AP So i redused the numerator as n'(n'+1)/2

OpenStudy (kanwal32):

@triciaal any ideas

OpenStudy (kanwal32):

@ParthKohli pls help

OpenStudy (triciaal):

sum of (n + 1)/n + 1 sorry i can't do more i have to go

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

@ParthKohli sir any ideas

Parth (parthkohli):

\[\sum \dfrac{1 + 2 + 3\cdots n}{n}\]\[= \sum \dfrac{n(n+1)/2}{n} \]\[= \sum \dfrac{n+1}{2}\]\[= \dfrac{1}{2}\sum(n+1)\]

OpenStudy (kanwal32):

thank u sir

Parth (parthkohli):

:) You have to do one more step of work though. Find \(\sum (n+1)\). It should be very easy.

OpenStudy (kanwal32):

yes sir

Parth (parthkohli):

Don't call me "sir" :P

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

thank u once again

Parth (parthkohli):

No problem :)